# Discrete-time thermodynamic speed limit

### v1

@@ Abstract
As a fundamental thermodynamic principle, speed limits reveal the lower bound of entropy production (EP) required for a system to transition from a given initial state to a final state. While various speed limits have been developed for continuous-time Markov processes, their application to discrete-time Markov chains remains unexplored. In this study, we investigate the speed limits in discrete-time Markov chains, focusing on two types of EP commonly used to measure the irreversibility of a discrete-time process: time-reversed EP and time-backward EP. We find that time-reversed EP satisfies the speed limit for the continuous-time Markov processes, whereas time-backward EP does not. Additionally, for time-reversed EP, we derive practical speed limits applicable to systems driven by cyclic protocols or with unidirectional transitions, where conventional speed limits become meaningless or invalid. We show that these relations also hold for continuous-time Markov processes by taking the time-continuum limit of our results. Finally, we validate our findings through several examples.
@@
In addition to the fluctuation theorems, numerous thermodynamic tradeoff relations have been discovered, surpassing the bound of the second law. These include the thermodynamic uncertainty relation (TUR) Barato and Seifert (2015); Hasegawa and Van Vu (2019); Horowitz and Gingrich (2020); Koyuk and Seifert (2020); Liu et al. (2020); Gupta and Busiello (2020); Van Vu and Hasegawa (2021); Pal et al. (2021a); Lee et al. (2021); Pal et al. (2021b); Park and Park (2021), entropic bounds Dechant and Sasa (2018), and fluctuation-response inequalities Dechant and Sasa (2020). These relations establish non-trivial lower bounds on EP in terms of generic non-equilibrium currents. They are instrumental in deriving power-efficiency tradeoff relations Shiraishi et al. (2016); Pietzonka and Seifert (2018); Lee et al. (2020); Oh and Baek (2023) and bounds on the turnover time scale of anomalous diffusions Hartich and Godec (2021). Moreover, these relations enable the inference of EP from experimental data without necessitating knowledge of the governing equations Li et al. (2019); Kim et al. (2022); Otsubo et al. (2022); Kim et al. (2022); Lee et al. (2023); Kwon and Baek (2023).
Another important finding is the discovery of thermodynamic speed limits. It establishes a bound on EP in terms of probability distance and the dynamical activity Shiraishi et al. (2018); Delvenne and Falasco (2021); Lee et al. (2022); Ito and Dechant (2020); Van Vu and Hasegawa (2021); Dechant (2022); Van Vu and Saito (2023). Recently, the speed limit has been utilized to derive the finite-time Landauer’s bound Delvenne and Falasco (2021); Lee et al. (2022); Van Vu and Saito (2023); Proesmans et al. (2020), while also interpreting it as part of a unified hierarchical tradeoff relations Kwon et al. (2023). Previous studies on speed limits Van Vu and Saito (2024) have focused on continuous-time Markov processes, where transition times are continuous. However, various interesting phenomena occur in discrete-time setups, such as Monte Carlo simulations, financial and stock price, Page Rank algorithm, and biochemical dynamics Van Kampen (1992). Despite their prevalence, the study of speed limits in discrete-time settings remains unexplored. Regarding the TURs, the conventional TUR derived for continuous-time Markov processes is violated in discrete-time processes, leading to a modification of the TUR formulation Proesmans and Van den Broeck (2017); Hasegawa and Van Vu (2019); Timpanaro et al. (2019); Liu et al. (2020). In this context, it is intriguing and important to investigate the effects of time discretization on speed limits.
@@
Here, we study speed limits for discrete-time Markov chains. For discrete-time processes, there are two widely accepted EP formulations: time-reversed EP and time-backward EP Igarashi (2022); Gaspard (2004); Proesmans et al. (2020); Lee (2018); Liu et al. (2020), which converge to the same value in continuous-time the limit. We investigate speed limits for these EPs and verify that the speed limit derived for continuous-time Markov processes still holds for time-reversed EP, while it is violated for time-backward EP. We quantify the degree of this violation using the Kullback-Leibler divergence between consecutive probability distributions. This indicates that the bound of the speed limit depends on the choice of EP formulation and that modification from the conventional form is required for time-backward EP. We also find the saturation conditions of the speed limits for time-reversed EP. In addition, we present practical speed limits for processes with unidirectional transitions or cyclic steady states, where conventional speed limits become invalid or meaningless. Our results, based on discrete-time processes, can be straightforwardly extended to continuous-time processes by taking their continuous-time limits. We also validate our findings through numerical calculations.
@@
This article is organized as follows. In Sec. II, we review the two formulations of EP for a discrete-time Markov chain. In Sec. III, we present the derivation of speed limits for these EPs in discrete-time Markov chains. Specifically, we derive the stepwise speed limit, which requires the full information of probability distributions at all time steps, and the section speed limit, which necessitates only partial information on probability distributions at limited time steps. The saturation condition for the speed limits is also discussed. In Sec. IV, we extend our result to a Markov chain with unidirectional transitions. In Sec. V, we numerically verify our results through three examples. Finally, we conclude in Sec. VI.
@@ II Entropy productions
Before discussing the speed limits, we first revisit the formulations of EP for a discrete-time Markov chain. We denote the $i$-th transition time occured during the process by $t_{i}$ for $i=0,1,...,N$ with a time gap $\Delta t=t_{i+1}-t_{i}$ and the probability of state $n$ at time $t_{i}$ by $P_{n}(t_{i})$. The Markov chain is characterized by the transition probability $M_{n,m}(t_{i})$, denoting the probability of transition from the state $m$ at $t_{i}$ to $n$ at $t_{i+1}$. We assume that the system obeys microscopic reversibility, i.e., $M_{n,m}(t)\neq 0$ if $M_{m,n}(t)\neq 0$. The case involving unidirectional transitions will be discussed in Sec. IV. In terms of $P_{n}(t_{i})$ and $M_{n,m}(t_{i})$, probability of state $n$ at $t_{i+1}$ is determined by the following equation:
@@
This equation is analogous to the following master equation for a continuous-time Markov process:
@@
where $R_{n,m}(t)$ is the transition rate at time $t$. EP rate for this continuous-time Markov process is given as Seifert (2012)
From Eq. (4), two formulations of EP for a discrete-time Markov chain can be deduced: (i) time-reversed EP $\tilde{\Sigma}$ Lee (2018); Liu et al. (2020) and (ii) time-backward EP $\Sigma$ Gaspard (2004); Proesmans et al. (2020). Integrating Eq. (4) over the infinitesimal time interval $dt=t_{i+1}-t_{i}$ leads to
[…]
where the system entropy change is defined as $\Delta S_{\textrm{sys}}(t_{i})\equiv\sum_{n}[P_{n}(t_{i})\ln P_{n}(t_{i})-P_{n}(t_{i+1})\ln P_{n}(t_{i+1})]$. Equation (3) is used to obtain Eq. (6) from Eq. (5). For an infinitesimal $dt$, Eq. (5) can be approximated as
$\Delta\Sigma_{\textrm{cont}}\approx\sum_{n,m}R_{n,m}(t_{i})dtP_{m}(t_{i})\ln\frac{R_{n,m}(t_{i})P_{m}(t_{i})}{R_{m,n}(t_{i})P_{n}(t_{i})}$
$=\sum_{n,m}M_{n,m}(t_{i})P_{m}(t_{i})\ln\frac{M_{n,m}(t_{i})P_{m}(t_{i})}{M_{m,n}(t_{i})P_{n}(t_{i})}\equiv\Delta\tilde{\Sigma}(t_{i}),$ (7)
where we used the relation $M_{n,m}(t_{i})=\delta_{n,m}+R_{n,m}(t_{i})dt+O(dt^{2})$. The formulation $\Delta\tilde{\Sigma}(t_{i})$ in Eq. (7) represents EP generated during one discrete-time step and has been used as a measure of irreversibility or EP in discrete-time processes. This EP is referred to as time-reversed EP. Note that the probabilities $P_{m}$ and $P_{n}$ in $\Delta\tilde{\Sigma}(t_{i})$ depend solely on a single time $t_{i}$, and thus, time-reversed EP shares the same mathematical form as the EP rate for continuous-time processes in Eq. (4). It is also interesting to note that the diagonal terms with $n=m$ in Eq. (7) do not contribute to $\Delta\tilde{\Sigma}(t_{i})$ as in continuous-time EP rate $\dot{\Sigma}_{\rm{cont}}(t)$.
Alternatively, starting from Eq. (6), one can derive another formulation of EP for discrete-time processes. Using the normalization condition $\sum_{n}M_{n,m}(t_{i})=1$ and Eq. (1), we can show
$\sum_{n,m}M_{n,m}(t_{i})P_{m}(t_{i})\ln\frac{P_{m}(t_{i})}{P_{n}(t_{i+1})}$
$=\sum_{n}P_{n}(t_{i})\ln P_{n}(t_{i})-\sum_{n}P_{n}(t_{i+1})\ln P_{n}(t_{i+1})$ (8)
which is identical to $\Delta S_{\rm sys}(t_{i})$. Plugging Eq. (II) into Eq. (6), we arrive at
$\Delta\Sigma_{\textrm{cont}}\approx\Delta S_{\textrm{sys}}(t_{i})+\sum_{n,m}R_{n,m}(t_{i})dtP_{m}(t_{i})\ln\frac{R_{n,m}(t_{i})}{R_{m,n}(t_{i})}$
$=\sum_{n,m}M_{n,m}(t_{i})P_{m}(t_{i})\ln\frac{M_{n,m}(t_{i})P_{m}(t_{i})}{M_{m,n}(t_{i})P_{n}(t_{i+1})}\equiv\Delta\Sigma(t_{i}).$ (9)
This formulation, $\Delta\Sigma(t_{i})$, is called time-backward EP. Different from time-reversed EP, $\Delta\Sigma(t_{i})$ necessitates the information of the next-time probability $P_{n}(t_{i+1})$ as well as current-time one $P_{n}(t_{i})$. The first line of Eq. (9) displays the decomposition of time-backward EP into the system entropy change and the contribution of the environment. In this context, time-backward EP is often employed in the studies related to thermodynamics. Unlike time-reversed EP, time-backward EP depends on the staying probability, meaning that events where the system remains in the same state also contribute to time-backward EP.
By using the log inequality $\ln(1/x)\geq 1-x$ for $x>0$, one can show that both EPs are always non-negative regardless of the magnitude of the time interval. In addition, $\Delta\tilde{\Sigma}\geq\Delta\Sigma$ can be proved by utilizing the same log inequality as follows:
@@
where $D[\mathbf{p}\parallel\mathbf{q}]=\sum_{n}p_{n}\ln(p_{n}/q_{n})$ denotes the Kullack-Leibler divergence of a probability $p_{n}$ with respect to a probability $q_{n}$. In the continuous-time limit, their difference $\Delta\Sigma^{\prime}$ vanishes at the order of the square of the time interval. Thus, both EPs become identical in the limit.
@@ III.1 Stepwise speed limit
$=A(t_{i})D[\mathbf{Q}(t_{i})\parallel\mathbf{Q}^{\textrm{T}}(t_{i})],$ (11)
where $Q_{n,m}(t_{i})$ is a joint probability defined as $Q_{n,m}(t_{i})\equiv M_{n,m}(t_{i})P_{m}(t_{i})/A(t_{i})$ with normalization condition $\sum_{n\neq m}Q_{n,m}(t_{i})=1$, $A(t_{i})\equiv\sum_{n\neq m}M_{m,n}(t_{i})P_{n}(t_{i})$ is the dynamical activity, and T denotes the transpose of a matrix, i.e., $Q^{\textrm{T}}_{n,m}(t_{i})=Q_{m,n}(t_{i})$.
The Kullback-Leibler divergence is lower bounded by a function of the total variation distance which is a measure of the distance between two distributions, defined as $d_{\textrm{TV}}[\mathbf{p},\mathbf{q}]\equiv\sum_{n}|p_{n}-q_{n}|/2$. This bound can be comprehensively expressed as Lee et al. (2022)
$D[\mathbf{Q}(t_{i})\parallel\mathbf{Q}^{\textrm{T}}(t_{i})]\geq h\left(d_{\textrm{TV}}[\mathbf{Q}(t_{i}),\mathbf{Q}^{\textrm{T}}(t_{i})]\right),$ (12)
where $h(x)$ represents a convex function satisfying Eq. (12). Several such functions have been found, for example, $h(x)=2x^{2}$ for Pinsker’s inequality Pinsker (1960) and $h(x)=\ln[(1+x)^{-1+x}/(1-x)]$ for Gilardoni’s inequality Gilardoni . Specifically, when the Kullback-Leibler divergence satisfies the symmetry condition $D[\mathbf{p}\parallel\mathbf{q}]=D[\mathbf{q}\parallel\mathbf{p}]$, the function $h(x)=x\ln[(1+x)/(1-x)]\equiv h_{\rm sym}(x)$ can be used and it provides the tightest bound Lee et al. (2022). As $D[\mathbf{Q}\parallel\mathbf{Q}^{\textrm{T}}]$ in Eq. (12) is symmetric, $h_{\rm sym}(x)$ can be employed. It is also worth noting that every $h(x)$ is a monotonically increasing function for $x\geq 0$.
Combining Eqs. (III.1) and (12), we can derive the lower bound of time-reversed EP in terms of dynamical activity $A(t_{i})$ and $\mathbf{Q}(t_{i})$ as follows:
@@
$\geq\left|\sum_{m(\neq n)}M_{n,m}(t_{i})P_{m}(t_{i})-M_{m,n}(t_{i})P_{n}(t_{i})\right|$ (14)
@@
$d_{\rm TV}[{\mathbf{Q}}(t_{i}),\mathbf{Q}^{\textrm{T}}(t_{i})]$ $=\sum_{n}\frac{\sum_{m(\neq n)}\left|Q_{n,m}(t_{i})-Q_{m,n}(t_{i})\right|}{2}$
@@
where $d_{i+1,i}\equiv d_{\rm TV}[\mathbf{P}(t_{i+1}),\mathbf{P}(t_{i})]$. Plugging Eq. (III.1) into Eq. (13) under the monotonically increasing property of $h(x)$ and summing over all $i$ yield
$\tilde{\Sigma}=\sum_{i=0}^{N-1}\Delta\tilde{\Sigma}(t_{i})\geq\sum_{i=0}^{N-1}A(t_{i})h\left\{\frac{d_{i+1,i}}{A(t_{i})}\right\}.$ (16)
@@
$\tilde{\Sigma}\geq$ $A_{\rm tot}h\left\{\sum_{i=0}^{N-1}\frac{d_{i+1,i}}{A_{\rm tot}}\right\}.$ (18)
This expression is applicable when $P(t_{0})=P(t_{N})$ and yields a nontrivial bound for this case. We note that $A_{\textrm{tot}}$ can be computed by averaging the number of transitions between all different states without the knowledge of $\mathbf{P}(t)$. On another note, a relation similar, in terms of Wasserstein distance, to Eq. (18) for continuous-time processes was reported in Ref. Van Vu and Saito (2024).
III.2 Section speed limit

### v2

# Speed Limits for Discrete-Time Markov Chains

@@ Abstract
As a fundamental thermodynamic principle, speed limits reveal the lower bound of entropy production (EP) required for a system to transition from a given initial state to a final state. While various speed limits have been developed for continuous-time Markov processes, their application to discrete-time Markov chains remains unexplored. In this study, we investigate the speed limits in discrete-time Markov chains, focusing on two types of EP commonly used to measure the irreversibility of a discrete-time process: time-reversed EP and time-backward EP. We find that time-reversed EP satisfies the speed limit for the continuous-time Markov processes, whereas time-backward EP does not. Additionally, for time-reversed EP, we derive practical speed limits applicable to systems driven by cyclic protocols or with unidirectional transitions, where conventional speed limits become meaningless or invalid. We show that these relations also hold for continuous-time Markov processes by taking the time-continuum limit of our results. Finally, we validate our findings through several examples: toy models, experimentally established cell-state dynamics, and theoretical computation.
@@
In addition to the fluctuation theorems, numerous thermodynamic tradeoff relations have been discovered, surpassing the bound of the second law. These include the thermodynamic uncertainty relation (TUR) Barato and Seifert (2015); Hasegawa and Van Vu (2019); Horowitz and Gingrich (2020); Koyuk and Seifert (2020); Liu et al. (2020); Van Vu and Hasegawa (2021); Pal et al. (2021a); Lee et al. (2021a); Pal et al. (2021b); Park and Park (2021), entropic bounds Dechant and Sasa (2018), and fluctuation-response inequalities Dechant and Sasa (2020); Euijoon et al. (2024). These relations establish non-trivial lower bounds on EP in terms of generic non-equilibrium currents. They are instrumental in deriving power-efficiency tradeoff relations Shiraishi et al. (2016); Pietzonka and Seifert (2018); Lee et al. (2020a); Oh and Baek (2023) and bounds on the turnover time scale of anomalous diffusions Hartich and Godec (2021). Moreover, these relations enable the inference of EP from experimental data without necessitating knowledge of the governing equations Li et al. (2019); Kim et al. (2022); Otsubo et al. (2022); Kim et al. (2022); Lee et al. (2023); Kwon and Baek (2024).
Another important finding is the discovery of thermodynamic speed limits Shiraishi et al. (2018); Gupta and Busiello (2020); Delvenne and Falasco (2024); Lee et al. (2022); Ito and Dechant (2020); Van Vu and Hasegawa (2021); Dechant (2022); Van Vu and Saito (2023). It establishes a bound on EP in terms of probability distance and the dynamical activity. Recently, the speed limit has been used to derive the finite-time Landauer’s bound Delvenne and Falasco (2024); Lee et al. (2022); Van Vu and Saito (2023); Proesmans et al. (2020) and interpret it within a unified hierarchy of tradeoff relations. Kwon et al. (2024).
Previous studies on speed limits have primarily focused on continuous-time Markov processes described by Master equations. However, discrete-time models are widely used across various research fields due to both practical and conceptual considerations. For example, the dynamics of interest may inherently evolve in discrete steps, as seen in generational updates in population Fisher (1999) or the evolution of binary states in computational algorithms Manzano et al. (2024). Even when the underlying processes are continuous, discrete-time models are often employed to analyze empirical data collected at fixed intervals, such as data from financial time series Bischi and Valori (2000), cell culture experiments Gupta et al. (2011), and climate time series Lee et al. (2024). Additionally, some coarse-grained discrete-time processes are used to infer physical quantities of the underlying continuous processes, such as a return map for the Lyapunov exponent Packard et al. (1980) and a milestoned process for entropy production Blom et al. (2024). Finally, some rule-based discrete-time models offer valuable theoretical insights with enhanced analytical tractability, contributing to the understanding of the routes to chaos May (1976), self-organized criticality Bak et al. (1988), and nonequilibrium phase transitions Achlioptas et al. (2009).
Despite the importance of discrete-time Markov processes, relatively few studies have addressed EP and thermodynamic relations in such systems. Some papers have defined EP in discrete-time Markov chains by relating it to the degree of irreversibility Gaspard (2004); Ge et al. (2006) and the violation of detailed balance, while others have addressed discrete-time EP as part of their primary goal of extending the TUR to discrete-time processes Shiraishi (2017); Proesmans and Van den Broeck (2017); Hasegawa and Van Vu (2019); Timpanaro et al. (2019); Liu et al. (2020). In particular, the violation of the TUR in discrete-time processes was first demonstrated in Shiraishi (2017), and explicit formulations were derived in subsequent studies Proesmans and Van den Broeck (2017); Hasegawa and Van Vu (2019); Timpanaro et al. (2019); Liu et al. (2020). However, the extensions of other thermodynamic relations, in particular speed limits, to discrete-time processes remain underexplored.
In addition, previous studies on speed limits have focused on systems without unidirectional transitions. However, systems with unidirectional transitions are often used to describe dynamics where reverse transitions are either extremely rare or missing due to incomplete experimental data Gupta et al. (2011); Baiesi et al. (2024). Physically relevant examples include totally asymmetric exclusion processes Derrida et al. (1993); Saha and Mukherji (2016); Soh et al. (2018), directed percolation Takeuchi et al. (2007), spontaneous emission Rahav and Harbola (2014), opinion dynamics Holley and Liggett (1975), stochastic resetting Evans and Majumdar (2011); Fuchs et al. (2016); Tal-Friedman et al. (2020); Chen et al. (2022); Mori et al. (2023), and theoretical computation Manzano et al. (2024), as well as various biological Gupta et al. (2011); Dharmaraja et al. (2015); Roldán et al. (2016); Bressloff (2020) and chemical processes Michaelis et al. (1913); Ting et al. (2009); Johnson and Goody (2011). While the TUR has been extended to systems with unidirectional transitions Pal et al. (2021a), a corresponding extension for speed limits remains unexplored. In this context, it is intriguing and important to investigate the effects of time discretization and unidirectional transitions on speed limits.
@@
Here, we study speed limits for discrete-time Markov chains. In the literature on discrete-time processes, at least two EP formulations can be found: time-reversed EP and time-backward EP Liu et al. (2020); Lee (2018); Gaspard (2004); Igarashi (2022); Proesmans and Van den Broeck (2017); Touchette (2009), both of which converge to the same value in the continuous-time limit. We investigate speed limits for these EPs and verify that the speed limit derived for continuous-time Markov processes still holds for time-reversed EP, while it is violated for time-backward EP. We quantify the degree of this violation using the Kullback-Leibler divergence between consecutive probability distributions. This indicates that the bound of the speed limit depends on the choice of EP formulation and that modification from the conventional form is required for time-backward EP. We also find the saturation conditions of the speed limits for time-reversed EP.
We extend the thermodynamic speed limit to two types of processes where the conventional speed limit fails or becomes trivial: (i) systems that return to their initial distribution Shiraishi et al. (2018); Lee et al. (2022) and (ii) systems with unidirectional transitions Manzano et al. (2024); Busiello et al. (2020); Rahav and Harbola (2014). In the first case, which is often used to describe thermal devices operating in a cyclic steady state Feldmann and Kosloff (2003); Park et al. (2019); Feldmann and Kosloff (2000); Lee et al. (2020b, 2021b); Chun and Park (2025), the probabilistic distance between the identical initial and final states is zero, causing the conventional speed limit bound to vanish and become meaningless. To address this limitation, we derive tighter bounds expressed in terms of either the sum of distances between consecutive states or the distance between two states at any intermediate time.
@@
This article is organized as follows. In Sec. II, we review the two formulations of EP for a discrete-time Markov chain. In Sec. III, we present the derivation of speed limits for these EPs in discrete-time Markov chains. Specifically, we derive the stepwise speed limit, which requires the full information of probability distributions at all time steps, and the interval speed limit, which necessitates only partial information on probability distributions at limited time steps. The saturation condition for the speed limits is also discussed. In Sec. IV, we extend our result to a Markov chain with unidirectional transitions. In Sec. V, we numerically verify our results through three examples. Finally, we conclude in Sec. VI.
@@ II Entropy productions
Before discussing the discrete-time speed limits, we first revisit the formulations of EP for a discrete-time Markov chain. A few references have explored the EP of a discrete-time Markov chain, with at least two types of EP being studied Liu et al. (2020); Lee (2018); Gaspard (2004); Igarashi (2022); Proesmans and Van den Broeck (2017); Touchette (2009). To explain these EPs, we introduce our notations for the discrete-time Markov chain.
We denote the time at which the $i$-th transition occurs as $t_{i}$ for $i=0,1,...,N$, with a time gap $\Delta t=t_{i+1}-t_{i}$, and the probability of being in state $n$ at time $t_{i}$ by $P_{n}(t_{i})$. The Markov chain is characterized by the transition probability $M_{n,m}(t_{i})$, denoting the conditional probability of transition from the state $m$ at $t_{i}$ to $n$ at $t_{i+1}$, $M_{n,m}(t)\equiv P(n,t_{i+1}|m,t_{i})$. We assume that every transition has a corresponding reverse transition, i.e., $M_{n,m}(t)\neq 0$ if $M_{m,n}(t)\neq 0$. The case involving unidirectional transitions will be discussed in Sec. IV.
In terms of $P_{n}(t_{i})$ and $M_{n,m}(t_{i})$, the probability of being in state $n$ at $t_{i+1}$ is given by the following equation:
@@
Equation (II) will be utilized to derive our discrete-time speed limits. This equation is analogous to the following master equation for a continuous-time Markov process:
@@
where $R_{n,m}(t)$ is the transition rate at time $t$. By dividing Eq. (II) by $\Delta t$ and taking the continuum limit, one can derive Eq. (3) from Eq. (II) when $M_{n,m}(t_{i})=\delta_{n,m}+R_{n,m}(t_{i})\Delta t+O(\Delta t^{2})$. The results of this paper can also be extended to deduce continuous-time results.
We note that, in general, time may not have a relevant physical meaning in discrete-time processes. For example, in computational algorithms, state transitions occur after each logical operation, without requiring a fixed time interval. In such cases, the discrete-time $t_{i}$ merely denote the index corresponding to the number of operations. For simplicity, we focus on cases where the time interval is fixed, while our results can be straightforwardly extended to general scenarios.
For discrete-time processes, two definitions of EP have been used in literature Igarashi (2022): (i) time-reversed EP $\tilde{\Sigma}$ Gaspard (2004); Touchette (2009); Proesmans and Van den Broeck (2017) and (ii) time-backward EP $\Sigma$ Lee (2018); Liu et al. (2020), which are given by
$\Delta\tilde{\Sigma}(t_{i})$ $\equiv\sum_{n,m}M_{n,m}(t_{i})P_{m}(t_{i})\ln\frac{M_{n,m}(t_{i})P_{m}(t_{i})}{M_{m,n}(t_{i})P_{n}(t_{i})},$ (4)
$\Delta\Sigma(t_{i})$ $\equiv\sum_{n,m}M_{n,m}(t_{i})P_{m}(t_{i})\ln\frac{M_{n,m}(t_{i})P_{m}(t_{i})}{M_{m,n}(t_{i})P_{n}(t_{i+1})}.$ (5)
Notably, their difference lies only in the time argument of $P_{n}(t)$. Thus, the time-reversed and time-backward EPs are identical when the system is in a steady-state $P_{n}(t)=P_{n}$.
By using the log inequality $\ln x+1-x\leq 0$ for $x>0$, one can show that both EPs are always non-negative regardless of the time interval. In addition, $\Delta\tilde{\Sigma}\geq\Delta\Sigma$ can be proved by using the same log inequality as follows:
@@
where $D[\mathbf{p}\parallel\mathbf{q}]=\sum_{n}p_{n}\ln(p_{n}/q_{n})$ denotes the Kullack-Leibler divergence of a probability $p_{n}$ with respect to a probability $q_{n}$. In the continuous-time limit, their difference, $\Delta\Sigma^{\prime}$, vanishes as the order of the square of the time interval. Thus, both EPs become identical in that limit.
Due to the difference in $P_{n}(t)$, the time-reversed and time-backward EP exhibit several significant differences. For instance, some properties of the EP rate in continuous-time processes are shared not by both, but solely by time-reversed EP. Their differences are explained in Appendix A.
@@ III.1 Stepwise speed limit
$=A(t_{i})D[\mathbf{Q}(t_{i})\parallel\mathbf{Q}^{\textsf{T}}(t_{i})],$ (7)
where $Q_{n,m}(t_{i})$ is a joint probability defined as $Q_{n,m}(t_{i})\equiv M_{n,m}(t_{i})P_{m}(t_{i})/A(t_{i})$ with normalization condition $\sum_{n\neq m}Q_{n,m}(t_{i})=1$, $A(t_{i})\equiv\sum_{n\neq m}M_{n,m}(t_{i})P_{m}(t_{i})$ is the dynamical activity, and T denotes the transpose of a matrix, i.e., $Q^{\textsf{T}}_{n,m}(t_{i})=Q_{m,n}(t_{i})$. The Kullback–Leibler divergence between two matrices is defined as
$D[\mathbf{A}\parallel\mathbf{B}]\equiv\sum_{n,m}A_{n,m}\ln\frac{A_{n,m}}{B_{n,m}}.$ (8)
Several inequalities establish the relation between the Kullback-Leibler divergence and the total variation distance, which is a measure of the distance between two distributions $\mathbf{p}$ and $\mathbf{q}$, defined as
$d_{\textrm{TV}}[\mathbf{p},\mathbf{q}]\equiv\frac{1}{2}\sum_{n}|p_{n}-q_{n}|.$ (9)
For example, Pinsker’s inequality Pinsker (1960) and Gilardoni’s inequality Gilardoni provide the bounds expressed as $D[\mathbf{p}\parallel\mathbf{q}]\geq 2(d_{\textrm{TV}}[\mathbf{p},\mathbf{q}])^{2}$ and $D[\mathbf{p}\parallel\mathbf{q}]\geq h_{\textrm{G}}\left(d_{\textrm{TV}}[\mathbf{p},\mathbf{q}]\right)$ with $h_{\textrm{G}}(x)=\ln[(1+x)^{-1+x}/(1-x)]$, respectively. Which of these two inequalities is tighter depends on the value of $d_{\textrm{TV}}[\mathbf{p},\mathbf{q}]$ Lee et al. (2022). When the Kullback-Leibler divergence is symmetric ($D[\bm{p}||\bm{q}]=D[\bm{q}||\bm{p}]$), a tighter bound can be derived Gilardoni (2006, 2008); Falasco et al. (2022). This bound can be obtained from the bound for the symmetrized Kullback–Leibler divergence, also known as the Jeffreys divergence $D_{\rm J}[\mathbf{p},\mathbf{q}]$ (See Appendix. B for its derivation):
$D_{\rm J}[\mathbf{p},\mathbf{q}]\equiv\frac{D[\mathbf{p}\parallel\mathbf{q}]+D[\mathbf{q}\parallel\mathbf{p}]}{2}\geq h\left(d_{\textrm{TV}}[\mathbf{p},\mathbf{q}]\right)$ (10)
where
$h(x)\equiv x\ln\frac{1+x}{1-x}.$ (11)
Thus, if Kullback-Leibler divergence is symmetric, $D_{\textrm{J}}[\mathbf{p},\mathbf{q}]=D[\mathbf{p}\parallel\mathbf{q}]$, we obtain Gilardoni (2006, 2008); Lee et al. (2022)
$D[\mathbf{p}\parallel\mathbf{q}]\geq h\left(d_{\textrm{TV}}[\mathbf{p},\mathbf{q}]\right).$ (12)
We note that $h(x)$ is convex and a monotonically increasing function for $x\geq 0$. Since
$D[\mathbf{Q}^{\textrm{T}}(t_{i})\parallel\mathbf{Q}(t_{i})]$ $=\sum_{n,m}Q_{m,n}(t_{i})\ln\frac{Q_{m,n}(t_{i})}{Q_{n,m}(t_{i})}$
$=\sum_{n,m}Q_{n,m}(t_{i})\ln\frac{Q_{n,m}(t_{i})}{Q_{m,n}(t_{i})}$
$=D[\mathbf{Q}(t_{i})\parallel\mathbf{Q}^{\textrm{T}}(t_{i})],$ (13)
the Kullback-Leibler divergence in Eq. (III.1) is symmetric. Consequently, we can apply Eq. (12) to Eq. (III.1), which yields
@@
$\geq\left|\sum_{m(\neq n)}\left[M_{n,m}(t_{i})P_{m}(t_{i})-M_{m,n}(t_{i})P_{n}(t_{i})\right]\right|$ (15)
@@
$d_{\rm TV}[{\mathbf{Q}}(t_{i}),\mathbf{Q}^{\textsf{T}}(t_{i})]$ $=\sum_{n}\frac{\sum_{m(\neq n)}\left|Q_{n,m}(t_{i})-Q_{m,n}(t_{i})\right|}{2}$
@@
where $d_{i+1,i}\equiv d_{\rm TV}[\mathbf{P}(t_{i+1}),\mathbf{P}(t_{i})]$. Plugging Eq. (III.1) into Eq. (14) under the monotonically increasing property of $h(x)$ and summing over all $i$ yields
$\tilde{\Sigma}\equiv\sum_{i=0}^{N-1}\Delta\tilde{\Sigma}(t_{i})\geq\sum_{i=0}^{N-1}A(t_{i})h\left(\frac{d_{i+1,i}}{A(t_{i})}\right).$ (17)
@@
$\tilde{\Sigma}\geq$ $A_{\rm tot}h\left(\sum_{i=0}^{N-1}\frac{d_{i+1,i}}{A_{\rm tot}}\right).$ (19)
This expression is applicable when $\mathbf{P}(t_{0})=\mathbf{P}(t_{N})$ and yields a nontrivial bound for this case. We note that $A_{\textrm{tot}}$ can be computed by averaging the number of transitions between all different states without the knowledge of $\mathbf{P}(t)$. On a related note, a similar relation in terms of the Wasserstein distance for continuous-time processes, analogous to Eq.(19), was reported in Ref.Van Vu and Saito (2024).
III.2 Interval speed limit